Find an equation of the line L that passes through the point (−2, 6)and satisfies the given condition. (Let x be the independent variable and y be the dependent variable.)
L is perpendicular to the line 4x + 3y = 6

Answer by Alan3354(69443)

Find an equation of the line L that passes through the point (-2,6)and satisfies the given condition. (Let x be the independent variable and y be the dependent variable.)
L is perpendicular to the line 4x + 3y = 6
----------------
Find the slope of the line 4x + 3y = 6
----
Lines perpendicular to it will have a slope that's the negative inverse.
----
Then it's y-y1 = m*(x-x1) where m = slope and (x1,y1) is the point (-2,6)